OpenStudy (anonymous):

In parallelogram DEFG, Dh=x+2, HF= 2y, GH=3x-3, and HE=5y+1. Find the values of x and y?

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

where is H ?

jhonyy9 (jhonyy9):

is the midle point of diagonales ?

OpenStudy (anonymous):

Yes

jhonyy9 (jhonyy9):

so than x+2=2y and 3x-3=5y+1 x+2=2y --- x=2y-2 3x-3=5y+1 3(2y-2)-3=5y+1 6y-6-3=5y+1 6y-9=5y+1 y=10 x=2(10)-2 x=20-2 x=18 x=18 y=10

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

Yes , thank you (:

jhonyy9 (jhonyy9):

good luck bye